OpenStudy (anonymous):

PLEASE HELP Solve 1n 4 + 1n (3x) = 2 Round Your Answer To The Nearest Hundredth

OpenStudy (anonymous):

Choices Are A. 1.13 B. 0.18 C. 1.41 D. 0.62

OpenStudy (butterflydreamer):

here are some log rules to help you out :) \[\ln a + \ln b = \ln ab\] \[\ln a = \log_{e} a\]

OpenStudy (butterflydreamer):

So first, using the log laws above. What would \[\ln 4 + \ln (3x) = ?\]

OpenStudy (butterflydreamer):

hint: 4 times 3x = ?

OpenStudy (anonymous):

Um, I'm not sure

OpenStudy (anonymous):

12x ?

OpenStudy (anonymous):

yeah you got \[\ln(12x)=2\] to solve

OpenStudy (anonymous):

I'm still confused?!

OpenStudy (butterflydreamer):

correct! :) So, \[\ln 4 + \ln 3x = 2 \] \[\ln 12x = 2\] use this rule to help you: \[\ln a =\log_{e} a\] and \[\log_{e} a = b \] \[e^{b}= a\]

OpenStudy (anonymous):

I'm still not sure how to do it ??? :-/

OpenStudy (anonymous):

12=e^2 ?

OpenStudy (butterflydreamer):

so first let's RE-WRITE "ln12x" in log form. \[\ln12x=\log_{e}12x \] So now we know: ln12x=2 \[\log_{e} 12x = 2\]

OpenStudy (butterflydreamer):

(sorry i made a typo before LOL) So, we want to use the rule to help solve for x: \[\log_{e} a = b\] where \[a = e^{b}\] apply this rule to \[\log_{e} 12x = 2\]

OpenStudy (butterflydreamer):

so, 12x = ?

OpenStudy (anonymous):

2 ?

OpenStudy (perl):

1n 4 + 1n (3x) = 2 by log rules: ln ( 4*3x) = 2 now raise both sides to the power of e

OpenStudy (perl):